i^41 (the answer in the back of the book is "i" -- I just can't figure out how that answer is arrived at)
Answer by scott8148(5891)

well...you can multiply i by itself 41 times, or...

looking at powers of i __ i^0=1, i^1=i, i^2=-1, i^3=-i, i^4=1, i^5=i...

the pattern shows a 4-step cycle __ so divide the exponent by 4 and use the remainder

i^41=i^(41-40)=i^1=i